OpenStudy (anonymous):

I got this one wrong to.. I tried changing the equation but it didn't fit right. A hospital sends an invoice to a patient. The patient schedules a payment plan in which she makes an initial payment of $1,451 the first month and then pays one-third of the previous month’s payment each of the following months until the invoice is paid in full. Approximately how much of the invoice will be paid after the fourth month?

OpenStudy (anonymous):

@Easyaspi314 I've been giving you medals to!

OpenStudy (anonymous):

So she has 3 payments to make, the 2nd, 3rd, and 4th payment. Each payment reduces the initial amount of $1,451 by 1/3. So this is an exponential decay problem. so it will be 1451(1-(1/3))^3 = 1451 (2/3)^3 = $429.93

OpenStudy (anonymous):

Exponential Decay: A(1 - r)^t A = initial mount = $14,51 in this problem r = fraction of percent t = time, in years, months, etc.. which is months in this case.

OpenStudy (anonymous):

@mathgeek0912

OpenStudy (anonymous):

A = $1,451 in this problem.